Kandon Enterprises, Inc., has two operating divisions; one manufactures machinery and the other breeds and sells horses. Both divisions are considered separate components as defined by generally accepted accounting principles. The horse division has been unprofitable, and, on November 15, 2021, Kandon adopted a formal plan to sell the division. The sale was completed on April 30, 2022. At December 31, 2021, the component was considered held for sale. On December 31, 2021, the company’s fiscal year-end, the book value of the assets of the horse division was $394,000. On that date, the fair value of the assets, less costs to sell, was $340,000. The before-tax loss from operations of the division for the year was $280,000. The company’s effective tax rate is 25%. The after-tax income from continuing operations for 2021 was $540,000.

Required: Prepare a partial income statement for 2021 beginning with income from continuing operations. Ignore EPS disclosures.

Prepare a partial income statement for 2021 beginning with income from continuing operations. Assume that the estimated net fair value of the horse division’s assets was $680,000, instead of $340,000. Ignore EPS disclosures.
Answer #1
1)
KEI
Partial Income Statement
For the year ended December 31st,2021
Income from Continuing Operations=(A) $   5,40,000.00
Discontinued Operations gain/loss
Loss from operations of discontinued component=(B) $ -3,34,000.00
Income tax benefit=(C )=(B)*25% $       83,500.00
Loss on discontinued operations=(D )=(B)+(C ) $ -2,50,500.00
Net Income=($540000-$250500) $   2,89,500.00
Loss from operations of discontinued component
Before tax loss from operation of the division $ -2,80,000.00
Impairment loss=(Fair Value< Book Value)=($394000-$340000) $     -54,000.00
Discontinued operations gain/loss $ -3,34,000.00
Income tax benefit($334000*25%) $       83,500.00
Net after tax loss on discontinued operations(-$334000+83500) $ -2,50,500.00
2)
KEI
Partial Income Statement
For the year ended December 31st,2021
Income from Continuing Operations=(A) $   5,40,000.00
Discontinued Operations gain/loss
Loss from operations of discontinued component=(B) $ -2,80,000.00
Income tax benefit=(C )=(B)*25% $       70,000.00
Loss on discontinued operations=(D )=(B)+(C ) $ -2,10,000.00
Net Income=($540000-$210000) $   3,30,000.00
Fair value> Book Value=No Impairment loss arises
$680000>$340000=No Impairment loss